# GS4: Generalizable Sparse Splatting Semantic SLAM

### v1

@@ Abstract
Traditional SLAM algorithms are excellent at camera tracking but might generate lower resolution and incomplete 3D maps. Recently, Gaussian Splatting (GS) approaches have emerged as an option for SLAM with accurate, dense 3D map building. However, existing GS-based SLAM methods rely on per-scene optimization which is time-consuming and does not generalize to diverse scenes well. In this work, we introduce the first generalizable GS-based semantic SLAM algorithm that incrementally builds and updates a 3D scene representation from an RGB-D video stream using a learned generalizable network. Our approach starts from an RGB-D image recognition backbone to predict the Gaussian parameters from every downsampled and backprojected image location. Additionally, we seamlessly integrate 3D semantic segmentation into our GS framework, bridging 3D mapping and recognition through a shared backbone. To correct localization drifting and floaters, we propose to optimize the GS for only 1 iteration following global localization. We demonstrate state-of-the-art semantic SLAM performance on the real-world benchmark ScanNet with an order of magnitude fewer Gaussians compared to other recent GS-based methods, and showcase our model’s generalization capability through zero-shot transfer to the NYUv2 and TUM RGB-D datasets.
@@ 1 Introduction
Semantic Visual Simultaneous Localization and Mapping (SLAM) is a long-standing challenge in computer vision that aims to reconstruct a dense 3D semantic map of an environment while simultaneously estimating camera poses from an input video. These dense semantic maps have widespread applications in autonomous driving, AR/VR and robot navigation, enabling robots to understand and interact with their surroundings more effectively by providing rich 3D spatial and semantic information.
@@
Traditional visual SLAM systems consist of several independent components, including keypoint detection, matching, and bundle adjustment [25]. Although traditional SLAM systems generally provide accurate localization, they struggle to generate accurate, dense 3D maps, limiting their effectiveness in robotics applications that require precise environmental interaction, such as manipulation tasks. Recent advances in differentiable rendering [23, 14] have opened up new opportunities for more streamlined yet more powerful visual SLAM approaches. Neural scene representations, such as Neural Radiance Fields (NeRF) [23], have been shown effective in generating accurate, dense 3D maps and have been successfully adopted as a 3D scene representation for visual SLAM [35, 45, 11]. However, volumetric rendering used in NeRF remains computationally expensive because a neural network needs to be trained on-the-fly, which leads to a compromise between reconstruction quality and training costs.
@@
Recently, 3D Gaussians have emerged as a promising 3D scene representation, offering fast, differentiable, and high quality rendering capabilities [15]. With these advantages, Gaussian-based representations have been shown to be well-suited for SLAM systems [13, 22]. However, existing approaches require test-time gradient-based optimization to estimate 3D Gaussians for each scene independently, which is computationally expensive and unsuitable for real-time applications. Moreover, these methods rely on heuristic-based Gaussian densification and pruning [15], often leading to dense Gaussian representations that do not scale well in large, real-world environments.
@@
In this paper, we propose a novel GS-based SLAM system GS4 (Generalizable Sparse Splatting Semantic SLAM), based on the idea of directly predicting 3D semantic Gaussians with a learned feed-forward network. With this approach, we no longer require the expensive test-time network learning or per-scene optimization. Besides, by incorporating an image recognition backbone, we can utilize the same features to predict both the Gaussian parameters as well as the semantics of each Gaussian, creating a unified approach for semantic SLAM without incorporating an external semantic segmentation algorithm.
GS4 directly predicts 3D semantic Gaussians from RGB-D images using a learned feed-forward network. Additionally, GS4 features a Gaussian refinement network that effectively merges Gaussians predicted from different video frames, replacing the handcrafted heuristics used for Gaussian densification and pruning in prior GS-based methods. Our refinement network enables a learned pruning strategy, producing a compact 3D Gaussian representation with an order of magnitude fewer Gaussians than competiting approaches. To adapt to global localization (bundle adjustment) schemes that would change past camera poses and point locations, we propose to optimize Gaussian parameters for 1 iteration after global localization. This allowed the algorithm to correct the Gaussian parameters while alleviating the common issue of “floaters” in feed-forward GS methods that float in the air and reduce map quality.
@@
We demonstrate that GS4 achieves state-of-the-art performance across all key metrics in localization, mapping, and segmentation on the real-world benchmark ScanNet, while using only of Gaussians compared to prior GS SLAM methods (Fig. 1). Furthermore, we highlight the generalization capability of our system via zero-shot transfer to the NYUv2 and TUM RGB-D datasets, which, to the best of our knowledge, is the first demonstration of semantic SLAM generalization in a modern neural SLAM system.
@@
We propose GS4, the first generalizable Gaussian splatting semantic SLAM approach on monocular RGB-D sequences. Results showed that GS4 obtains state-of-the-art on real ScanNet scenes and also zero-shot generalizes to the real NYUv2 and TUM RGB-D datasets without any fine-tuning.
Our proposed Gaussian refinement network effectively merges Gaussians from different frames into a 3D representation, while significantly reducing the number of Gaussians required to represent a scene.
Our proposed 1-iteration Gaussian optimization significantly improves reconstruction quality with a small additional computational cost.
@@ 2 Related Work
Traditional SLAM: Early visual SLAM methods [25] demonstrated robust localization through effective keypoint detection and matching, which resulted in sparse 3D reconstructions. While these approaches provided reliable localization, the sparse nature of the reconstructed maps limited their utility in applications requiring detailed 3D maps. To address this issue, dense visual SLAM [16, 6] focused on constructing detailed maps to support applications like augmented reality (AR) and robotics. Prior methods [2, 8, 27, 1, 38, 29] achieved dense mapping by employing representations based on Signed Distance Fields (SDF), rather than relying on sparse representations such as point clouds or grids. However, these approaches often suffer from over-smoothed reconstruction, failing to capture fine details crucial for certain tasks.
@@
NeRF-based SLAM: Neural Radiance Fields (NeRF) [23] has gained popularity as a 3D scene representation due to its ability to generate accurate and dense reconstructions. NeRF employs Multi-Layer Perceptron (MLP) to encode scene information and performs volume rendering by querying opacity and color along pixel rays. Methods such as iMAP [35], NICE-SLAM [45], and ESLAM [11] incorporate this implicit scene representation into SLAM, leveraging NeRF’s high-fidelity reconstructions to improve both localization and mapping. However, the volumentric rendering process in NeRF is computationally expensive, often requiring trade-offs such as limiting the number of pixels during rendering, These trade-offs, while improving efficiency, may compromise the system’s accuracy in both localization and mapping.
@@
GS-based SLAM: 3D Gaussian Splatting (3DGS) [15] employs splatting rasterization instead of ray marching. This approach iterates over 3D Gaussian primitives rather than marching along rays, resulting in a more expressive and efficient representation capable of capturing high-fidelity 3D scenes with significantly faster rendering times. With the efficiency and rapid rasterization capabilities of 3D Gaussian Splatting, GS-based SLAM systems achieve improved accuracy and speed in dense scene reconstruction. SplaTAM [12] introduces silhouette-guided rendering to facilitate structured map expansion, enabling efficient dense visual SLAM. Gaussian Splatting SLAM [22] integrates novel Gaussian insertion and pruning techniques. Additionally, GS-ICP SLAM [9] combines ICP with 3DGS, achieving both faster computational speed and superior map quality. Expanding upon these advancements, SGS-SLAM [18] extends 3D Gaussian representations to include semantic scene understanding. However, existing GS-based SLAM methods employ per-scene optimization, requiring iterative refinement of Gaussians initialized from keyframes through rendering supervision. As a result, SGS-SLAM relies on additional segmentation models to predict semantic labels for each image, creating computational overhead.
@@
Feed-forward Models for GS: Recent research has introduced feed-forward approaches for scene-level 3DGS reconstruction using generalizable models [3, 4, 19]. Unlike previous methods, GS-LRM [40] avoids specialized 3D structural designs, instead uses an attention-based transformer to achieve state-of-the-art results. However, to the best of our knowledge, feed-forward models have been applied on a small number of images and have not yet been introduced in GS-based semantic SLAM approaches.
@@ 3.1 Gaussian Splatting
We represent a 3D map using a set of anisotropic 3D Gaussians used in Gaussian Splatting. Each Gaussian is characterized by RGB color , center position , scale , quaternion , opacity and semantic class vector , where N is the number of classes.
@@
where is a quantity of a pixel to be rendered, which can be color, depth or semantic label, and is that quantity of the -th 3D Gaussian, while
defines the visibility of the -th splatted Gaussian. Here, denotes the offset between the pixel coordinates and the 2D mean of the -th Gaussian. Following [13], We also render a silhouette image to determine visibility:
@@ 3.2 Gaussian Prediction and Refinement
Our proposed Gaussian prediction network takes RGB-D images as input and predicts 3D Gaussian parameters. These parameters also include semantic labels, enabling the rendering of photometric, geometric, and semantic views. Next, the Gaussian refinement network processes Gaussians predicted from different video frames and learns to merge them, constructing a consistent 3D scene representation.
@@ 3.2.1 Backbone for Gaussian Prediction
We train a transformer model to regress 3D GS parameters from an image with a known camera pose (from tracking, described in Sec. 3.3.1), while simultaneously assigning semantic labels to these 3D Gaussians. We start with ODIN [10], a model that starts from multi-view RGB-D images with a transformer architecture that alternates between 2D within-view and 3D cross-view attention for information fusion over multiple views. To further enhance efficiency, we have made the following modifications to the ODIN architecture:
We modify the ODIN architecture by removing the 3D attention layers between all stages of the 2D backbone, retaining only the 3D attention layers after their multi-scale deformable attention layer. This adjustment reduces computational costs while maintaining performance.
ODIN originally utilizes voxel-based operations, such as mean pooling of features and XYZ coordinates within occupied voxels in 3D space, which are beneficial for dense point clouds but can be computationally intensive. Given that our inference involves single-view input (i.e., sparser point cloud) at a time, we eliminate these voxel-based operations, resulting in faster processing without compromising performance.
ODIN utilizes 3D interpolation to upsample low-resolution feature maps. We replace this step with the point-based upsampling layer from AutoFocusFormer [46], which is more efficient.
These modifications collectively enhance the efficiency of the ODIN architecture ( speedup compared to the original), making it more suitable for our SLAM system.
Generating Prediction Tokens with ODIN. We concatenate an RGB image and a depth image resulting in a 4-channel feature map that is fed into the model:
where denotes the total number of prediction tokens per image. The variable represents the network level at which Gaussians are predicted. The backbone consists four levels, each processing a progressively downsampled set of tokens. If the second level is chosen, the feature map has a spatial resolution of , resulting in tokens per image. Here, represents the encoder features of an -th token from the backbone, while denotes the corresponding decoder features at the same resolution.
Processing Prediction Tokens with Transformer. Given the selected prediction level, we concatenate the encoder features and decoder features for each token , and process them through local-attention transformer layers [46]: to obtain the final features for the -th token, integrating information from both the encoder and decoder.
@@
Here, and represent the offsets for the 2D position (, ) of Gaussian in image space, while is the offset for the noisy depth obtained from the depth image. These offsets are added to the original values, which are then back-projected into 3D space using the camera’s intrinsic and extrinsic parameters, yielding the 3D center position . Similarly, represents the offset for the RGB value, which is added to the original RGB value observed from the image, yielding the final RGB value for each Gaussian.
Besides Gaussian parameters, the mask decoder head in ODIN predicts segmentation labels for the input image. By indexing these labels based on the tokens’ positions, we assign a semantic class vector to each Gaussian. Finally, each Gaussian is also assigned the final feature vector of its corresponding token for use in the subsequent Gaussian refinement stage.
To supervise the prediction of semantic segmentation, we follow the setup in Mask2Former [5]. We denote the corresponding segmentation loss by . In addition, we render images at supervision views—comprising the current input view and randomly selected novel views that overlap with the current input—using the predicted Gaussians from the current input, and minimize RGB-D and semantic rendering loss. For novel view supervision, we focus solely on areas visible in the input view, ensuring that the optimization process focuses on regions consistently observed across both input and novel views. We explain the loss functions used during training below.
RGB Rendering Loss. Following previous work [40, 47], we use a combination of the Mean Squared Error (MSE) loss and Perceptual loss:
where is the weight for the perceptual loss.
@@
Semantic Rendering Loss. For semantic rendering, we use the cross entropy loss:
where the rendered semantic image has N channels, each corresponding to a different semantic category.
Overall Training Loss. Our total loss comprises multiple rendering losses and the segmentation loss :
where we use , and .
@@ 3.2.2 Gaussian Refinement Network
In our SLAM system, we merge predicted Gaussians from a new frame into the 3D map. To ensure that the combined set of Gaussians accurately represents the scene, we propose a novel Gaussian Refinement Network to refine both the existing Gaussians in the 3D map and the newly added ones, enabling their effective merging. The input to the network includes the features and 3D positions of all Gaussians from both the 3D map and the new frame. We process these using several local-attention transformer layers with 3D neighborhoods to fuse and update the features for each Gaussian. Subsequently, MLP layers predict updates , , and for each Gaussian. These updates refine the Gaussians to accurately render both current and previous views. To supervise the network, we render the current input view along with overlapping previous views. The total training loss is defined as follows:
where we use , and .
During testing time, we introduce a threshold to manage the uncertainty of each Gaussian. Once a Gaussian has been updated times by the refinement network, we consider its uncertainty sufficiently reduced and exclude them from further updates. We set in our experiments. After Gaussian refinement, we prune Gaussians whose updated opacity falls below 0.005, effectively removing those that have become unimportant after merging.

### v2

@@ Abstract
Traditional SLAM algorithms excel at camera tracking, but typically produce incomplete and low-resolution maps that are not tightly integrated with semantics prediction. Recent work integrates Gaussian Splatting (GS) into SLAM to enable dense, photorealistic 3D mapping, yet existing GS-based SLAM methods require per-scene optimization that is slow and consumes an excessive number of Gaussians. We present GS4, the first generalizable GS-based semantic SLAM system. Compared with prior approaches, GS4 runs 10× faster, uses 10× fewer Gaussians, and achieves state-of-the-art performance across color, depth, semantic mapping and camera tracking. From an RGB-D video stream, GS4 incrementally builds and updates a set of 3D Gaussians using a feed-forward network. First, the Gaussian Prediction Model estimates a sparse set of Gaussian parameters from input frame, which integrates both color and semantic prediction with the same backbone. Then, the Gaussian Refinement Network merges new Gaussians with the existing set while avoiding redundancy. Finally, we propose to optimize GS for only 1-5 iterations that corrects drift and floaters when significant pose changes are detected. Experiments on the real-world ScanNet and ScanNet++ benchmarks demonstrate state-of-the-art semantic SLAM performance, with strong generalization capability shown through zero-shot transfer to the NYUv2 and TUM RGB-D datasets.
@@ 1 Introduction
Simultaneous Localization and Mapping (SLAM) is a long-standing challenge in computer vision, aiming to reconstruct a 3D map of an environment while simultaneously estimating camera poses from a video stream. Semantic visual SLAM extends this goal by producing dense maps enriched with semantic labels, enabling applications in autonomous driving, AR/VR, and robotics. By combining geometric reconstruction with object-level understanding, semantic SLAM provides rich 3D spatial and semantic information that allows robots and other systems to navigate and interact with their surroundings more effectively.
@@
Traditional visual SLAM systems consist of several independent components, including keypoint detection, feature matching, and bundle adjustment (Mur-Artal et al., 2015; Mur-Artal & Tardós, 2017; Campos et al., 2021). Their scene representations are typically low-resolution voxels, which limit geometric detail. Thus, although these systems generally provide accurate camera localization, they struggle to generate dense, high-quality 3D maps, which are required for robotics applications such as mobile manipulation. Recent advances in differentiable rendering (Mildenhall et al., 2020; Kerbl et al., 2023a) introduces new options for scene representation in visual SLAM. For example, neural scene representations such as Neural Radiance Fields (NeRF) (Mildenhall et al., 2020) have been successfully adopted in SLAM frameworks (Sucar et al., 2021; Zhu et al., 2022b; Johari et al., 2023); however, NeRF requires hours of per-scene optimization, making it computationally expensive and forcing a trade-off between reconstruction quality and training cost.
@@
Recently, 3D Gaussians have emerged as a powerful 3D scene representation, offering fast, differentiable, and high-quality rendering capabilities (Kerbl et al., 2023b). Leveraging these advantages, Gaussian-based representations have proven highly effective for SLAM systems (Keetha et al., 2024; Matsuki et al., 2024). However, existing approaches still rely on test-time, gradient-based optimization to estimate 3D Gaussians for each scene independently, which is computationally expensive and unsuitable for real-time applications. In addition, these methods depend on heuristic Gaussian densification and pruning strategies (Kerbl et al., 2023b), often producing overly dense representations that fail to scale to large, real-world environments.
@@
In this paper, we propose GS4 (Generalizable Sparse Splatting Semantic SLAM), the first generalizable Gaussian-splatting–based SLAM system, which directly predicts 3D semantic Gaussians using a learned feed-forward network, eliminating the need for expensive per-scene optimization. By integrating an image recognition backbone, GS4 jointly reconstructs geometry, color, and semantic labels of the environment without relying on any external semantic-segmentation modules.
GS4 begins with the Gaussian Prediction Model that infers a sparse set of 3D semantic Gaussians from each incoming RGB-D frame in a feed-forward manner. Next, the Gaussian Refinement Network integrates these newly predicted Gaussians with the evolving 3D map, replacing the handcrafted heuristics traditionally used for Gaussian densification and pruning. This learned refinement strategy yields a compact representation with an order-of-magnitude fewer Gaussians than competing methods. Finally, after the global localization (bundle adjustment) step from the camera tracking module updates camera poses and Gaussian locations, we perform a lightweight few-iteration (only 15) optimization of Gaussian parameters to preserve rendering fidelity and mitigate the “floater” artifacts common in feed-forward GS approaches.
@@
We demonstrate that GS4 achieves state-of-the-art performance across all key metrics in localization, mapping, and segmentation on the real-world benchmark ScanNet (Fig. 2), while using only of Gaussians compared to prior GS SLAM methods (Fig. 2). Furthermore, we highlight the generalization capability of our system via zero-shot transfer to the NYUv2 and TUM RGB-D datasets, which, to the best of our knowledge, is the first demonstration of zero-shot semantic SLAM generalization in a modern neural SLAM system.
@@
We propose GS4, the first generalizable Gaussian splatting semantic SLAM approach on monocular RGB-D sequences. Results showed that GS4 obtains state-of-the-art on real ScanNet and ScanNet++ scenes, and also zero-shot generalizes to the real NYUv2 and TUM RGB-D datasets without any fine-tuning.
Our proposed Gaussian refinement network effectively merges Gaussians from different frames into a 3D representation, while significantly reducing the number of Gaussians required to represent a scene to only of prior work.
Our proposed few-iteration Gaussian optimization significantly improves reconstruction quality with a small additional computational cost.
@@ 2 Related Work
Traditional SLAM: Early visual SLAM methods (Mur-Artal et al., 2015) demonstrated robust localization through effective keypoint detection and matching, which resulted in sparse 3D reconstructions. While these approaches provided reliable localization, the sparse nature of the reconstructed maps limited their utility in applications requiring detailed 3D maps. To address this issue, dense visual SLAM (Kerl et al., 2013; Czarnowski et al., 2020) focused on constructing detailed maps to support applications like augmented reality (AR) and robotics. Prior methods (Canelhas et al., 2013; Dai et al., 2017b; Newcombe et al., 2011; Bylow et al., 2013; Whelan et al., 2013; Prisacariu et al., 2017) employ representations based on Signed Distance Fields (SDF), rather than relying on sparse representations such as point clouds or grids. However, these approaches often suffer from over-smoothed reconstruction, failing to capture fine details crucial for certain tasks.
@@
NeRF-based SLAM: Neural Radiance Fields (NeRF) (Mildenhall et al., 2020) gained popularity as a 3D scene representation due to its ability to generate accurate and dense reconstructions. NeRF employs Multi-Layer Perceptron (MLP) to encode scene information and performs volume rendering by querying opacity and color along pixel rays. Methods such as iMAP (Sucar et al., 2021), NICE-SLAM (Zhu et al., 2022b), and ESLAM (Johari et al., 2023) incorporate this implicit scene representation into SLAM, leveraging NeRF’s high-fidelity reconstructions to improve both localization and mapping. DNS-SLAM (Li et al., 2023) further incorporates semantic information into the framework. However, the volumentric rendering process in NeRF is costly, often requiring trade-offs such as limiting the number of pixels during rendering, These trade-offs, while improving efficiency, may compromise the system’s accuracy in both localization and mapping.
@@
GS-based SLAM: 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023b) employs splatting rasterization instead of ray marching. This approach iterates over 3D Gaussian primitives rather than marching along rays, resulting in a more expressive and efficient representation capable of capturing high-fidelity 3D scenes with significantly faster rendering speed. Hence, GS-based SLAM systems achieve improved accuracy and speed in dense scene reconstruction. SplaTAM (Keetha et al., 2023) introduces silhouette-guided rendering to support structured map expansion, enabling efficient dense visual SLAM. Gaussian Splatting SLAM (Matsuki et al., 2024) integrates novel Gaussian insertion and pruning strategies, while GS-ICP SLAM (Ha et al., 2024) and RTG-SLAM (Peng et al., 2024) combine ICP with 3DGS to achieve both higher speed and superior map quality. Expanding upon these advancements, SGS-SLAM (Li et al., 2024b), OVO-SLAM (Martins et al., 2024), SemGauss-SLAM (Zhu et al., 2025) and GS3LAM (Li et al., 2024a) extend 3D Gaussian representations to include semantic scene understanding. However, existing GS-based SLAM methods employ per-scene optimization, requiring iterative refinement of Gaussians initialized from keyframes through rendering supervision. As a result, they all rely on additional segmentation models to predict semantic labels for each image, creating computational overhead.
@@
Feed-forward Models for GS: Recent research has introduced feed-forward approaches for scene-level 3DGS reconstruction using generalizable models (Charatan et al., 2024; Chen et al., 2024; Liu et al., 2025). Unlike previous methods, GS-LRM (Zhang et al., 2024) avoids specialized 3D structural designs, instead using a transformer to achieve state-of-the-art results. However, to the best of our knowledge, feed-forward models have been applied on a small number of images and have not yet been introduced in GS-based semantic SLAM approaches with thousands of frames.
@@ 3.1 Gaussian Splatting
We represent a 3D map using a set of anisotropic 3D Gaussians. Each Gaussian is characterized by RGB color , center position , scale , quaternion , opacity and semantic class vector , where N is the number of classes.
@@
where is a quantity of a pixel to be rendered, which can be color, depth or semantic label, and is that quantity of the -th 3D Gaussian, while is its visibility, computed from opacity and covariance parameters (determined by rotation and scale). Following Keetha et al. (2024), We also render a silhouette image to determine visibility .
@@ 3.2 Gaussian Prediction and Refinement
Our proposed Gaussian prediction network (Fig. 3) takes RGB-D images as input and predicts 3D Gaussian parameters. Importantly, the backbone generates features that can predict semantic labels (e.g. trained from 2D segmentation tasks), enabling the rendering of photometric, geometric, and semantic views. Next, the Gaussian refinement network processes Gaussians predicted from a new frame and learns to merge them with the 3D scene representation computed from prior frames.
@@ 3.2.1 Backbone for Gaussian Prediction
We train a transformer model to regress 3D GS parameters from an image with a known camera pose (from tracking, described in Sec. 3.3.1), while simultaneously assigning semantic labels to these 3D Gaussians. We start with a pre-trained 2D image segmentation model such as Mask2Former (Cheng et al., 2022) or AutoFocusFormer (Ziwen et al., 2023), which encodes an image into encoder tokens and decoder tokens (from their image decoder) at several progressively downsampled levels , with usually. We concatenate an RGB image and a depth image resulting in a 4-channel feature map that is fed into the model:
where denotes the total number of prediction tokens per image. The variable represents the network level at which Gaussians are predicted. If the second level is chosen, the feature map usually has a spatial resolution of , resulting in tokens per image.
Processing Prediction Tokens with Transformer. Given the selected prediction level, we concatenate the encoder features and decoder features for each token , and process the resulting tokens using local-attention transformer layers in the image space to obtain the final features for the -th token, integrating information from both the encoder and decoder.
@@
Here, and represent the offsets from the 2D position (, ) of the token in the image space, while is the offset for the noisy depth obtained from the depth image. These offsets are added to the original values, which are then backprojected into 3D space using the intrinsic and extrinsic parameters of the camera, yielding the 3D center position . Similarly, represents the offset for the RGB values, obtained from the downsampled image, where each token corresponds to a single pixel. Adding the offset to this value yields the final RGB color for each Gaussian. Besides Gaussian parameters, the mask decoder head in the segmentation model predicts token-level semantic segmentation label vector for the input image, which we then assign to the associated Gaussian. Finally, each Gaussian is assigned the final feature vector of its corresponding token for the subsequent Gaussian refinement stage.
To supervise the prediction of semantic segmentation, we follow the setup in Mask2Former (Cheng et al., 2022). We denote the corresponding segmentation loss by . In addition, we render images at supervision views—comprising the current input view and randomly selected novel views that overlap with the current input—using the predicted Gaussians from the current input, and minimize RGB-D and semantic rendering loss. For novel view supervision, we focus solely on areas visible in the input view, ensuring that the optimization process focuses on regions consistently observed across both input and novel views. We explain the loss functions used during training below.
RGB Rendering Loss. Following previous work Zhang et al. (2024); Ziwen et al. (2024), we use a combination of the Mean Squared Error (MSE) loss and Perceptual loss: where is the weight for the perceptual loss.
@@
Semantic Rendering Loss. For semantic rendering, we use the cross entropy loss: where the rendered semantic image has N channels, each corresponding to a different semantic category.
Overall Training Loss. Our total loss comprises multiple rendering losses and the segmentation loss : , where we use , and .
@@ 3.2.2 Gaussian Refinement Network
The previous subsection predicts Gaussian parameters from a single frame. In our SLAM system, as new frames arrive, we insert Gaussians from the frame into unmapped regions of the current 3D reconstruction. We perform co-visibility, which involves rendering a silhouette image for the new frame to identify the regions where new Gaussians should be inserted. To ensure that the combined set of Gaussians accurately represents the scene, we propose a novel Gaussian Refinement Network to refine both the existing Gaussians in the 3D map and the newly added ones, enabling their effective merging. The input to the network includes the features and 3D positions of all Gaussians from the 3D map that are visible in the new frame, as well as Gaussians from the new frame. We process these using several local-attention transformer layers with 3D neighborhoods in the world coordinate system to fuse and update the features for each Gaussian. Subsequently, MLP layers predict updates , , and for each Gaussian. These updates refine the Gaussians to accurately render both current and previous views. To supervise the network, we render the current view along with previous overlapping views. The total training loss is:
where we use , and . After Gaussian refinement, we prune Gaussians whose updated opacity falls below 0.005, effectively removing those that have become unimportant after merging. These merging-pruning steps lead to a significantly reduced number of Gaussians in the final 3D map with little performance impact.
During testing time, we introduce a threshold to manage the uncertainty of each Gaussian. Once a Gaussian has been updated times by the refinement network, we consider its uncertainty sufficiently reduced and exclude them from further updates. We set in our experiments.